Three identical conductive balls of equal size and shape, labeled A, B, and C carry a charge qA=+3C, qB=+2C, and qC= -4C, respectively. The following three steps are carried out sequentially:

1. A is brought to touch B and the separated.
2. A and C touch each other and separated.
3. B touches C and separated.

What is the magnitude of the charge, in COULOMBS, on ball B?
Answer #1

charge on B and A after contact

qa = qb = (3+2)/2 = 2.5 C

charge on C after touch with qa

qc = qa = ( 2.5 - 4) / 2 = - 0.75 C

net charge on B

qb' = (2.5- 0.75) /2 = 0.875 uC

========

Comment before rate in case any doubt, will reply for sure,, goodluck